OpenStudy (anonymous):

"(Adding & Subtracting Rational Expressions) The distance between Atlanta, Georgia, and Albuquerque, Mexico, is about 1270 miles. The ground speed for jet traffic from Albuquerque to Atlanta is 18% faster than the ground speed from Atlanta to Albuquerque in miles per hour. What is the simplified expression for the round trip flying time." Can someone help me find the answer to this or at least come up with the rational expression? Thank you.

OpenStudy (anonymous):

average speed = total distance / total time Let t = time from ATL to ABQ speed ATL→ABQ = 1270/t speed ABQ→ATL = 1.18(speed ATL→ABQ) = 1.18(1270/t) = (59/50)(1270/t) = 7493/(5t) Add them together and divide by 2 to find the average speed. Total distance = 2540 total time = t + 59/(50t)

OpenStudy (anonymous):